### v1

# Conformal Validity Guarantees Exist for Any Data Distribution

Abstract
As machine learning (ML) gains widespread adoption, practitioners are increasingly seeking means to quantify and control the risk these systems incur. This challenge is especially salient when ML systems have autonomy to collect their own data, such as in black-box optimization and active learning, where their actions induce sequential feedback-loop shifts in the data distribution. Conformal prediction has emerged as a promising approach to uncertainty and risk quantification, but existing variants either fail to accommodate sequences of data-dependent shifts, or do not fully exploit the fact that agent-induced shift is under our control. In this work we prove that conformal prediction can theoretically be extended to any joint data distribution, not just exchangeable or quasi-exchangeable ones, although it is exceedingly impractical to compute in the most general case. For practical applications, we outline a procedure for deriving specific conformal algorithms for any data distribution, and we use this procedure to derive tractable algorithms for a series of agent-induced covariate shifts. We evaluate the proposed algorithms empirically on synthetic black-box optimization and active learning tasks.
Machine Learning, ICML, conformal prediction, AI agents, exchangeability, covariate shift, active learning, black-box optimization
1 Introduction
Quantifying the uncertainty of predictions from artificial intelligence (AI) and machine learning (ML) models is often imperative to managing the risks associated with their downstream use. Principled uncertainty quantification (UQ) is thus especially crucial in real-world scenarios where AI/ML systems are used to inform high-stakes decisions, or where they may even act autonomously. Unfortunately, it is precisely when uncertainty estimation is most vital when it is often most difficult to quantify reliably. For a prime example, consider the increasingly common setting where AI systems are able to take actions of their own, thereby transforming them from passive observers into active agents: common instances include active learning, black-box optimization, and reinforcement learning. In all these cases, merely enabling an AI model to explore by actively selecting the next datapoint induces feedback-loop shifts in the data distribution, which can accumulate over time and cause standard UQ methods to deteriorate or break down entirely.
Conformal prediction (CP) (Vovk et al., 2005) is an increasingly popular UQ framework because it provides coverage guarantees without knowledge of the exact distributional form of the data. Even so, standard CP methods do make one very strong assumption about the data distribution: namely, that the data are exchangeable (IID is a special case). Informally, for time-series data, exchangeability means the distribution is “time invariant”; meanwhile in a batch setting, it means the distribution remains constant between training and test batches. Exchangeability thus excludes many real-world settings such as dynamic time series or data generated from (or queried by) active AI agents, which can exhibit stark time-dependent or batch-dependent distribution shifts.
This work builds on that of Tibshirani et al. (2019), who introduced a weighted generalization of CP where the traditional exchangeability requirement is relaxed to a condition called weighted exchangeability. For instance, data that are independent but not identically distributed are weighted exchangeable. However, the conventional wisdom is still that conformal prediction methods require some notion of “quasi-exchangeability” to achieve a valid coverage guarantee. [^1] For instance, Fannjiang et al. (2022) introduced a similar assumption called “pseudo-exchangeability” to study data shifts with limited dependencies of the test point on the observed data. We challenge this common perception via the following theoretical and practical contributions:
  • Our main theoretical contribution is to present a general view of (weighted) conformal prediction, drawing heavily upon the analysis of Tibshirani et al. (2019). In particular, we prove that conformal prediction can in theory be extended to any joint distribution of data with a valid probability density function $f$, including non-exchangeable ones with sequential dependencies, although the fully general form is not practically feasible to compute (factorial runtime and requiring knowledge of $f$). From this general perspective, “quasi-exchangeability” definitions (including exchangeability and weighted exchangeability) are no longer necessary conditions for conformal prediction theory; instead, from this view, these assumptions’ only role is to specify conditions for practical tractability.
  • For practical applications, we present a general procedure for deriving specific weighted CP algorithms for any given data distribution, and we demonstrate how this procedure can be used to derive weighted CP methods with valid coverage guarantees for sequential feedback-loop covariate shifts that characterize common AI/ML agent scenarios. We moreover show that these methods can be realized in practice via empirical evaluation on in silico protein design and active learning tasks, which verify that we maintain coverage when state-of-the-art baselines fail.
2 Background
Assume we are given an initial dataset $Z_{1:n}:=\{Z_{i}\}_{i=1}^{n}$, where $Z_{i}:=(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y}$ are real-valued ($\mathcal{X}\times\mathcal{Y}\subseteq\mathbb{R}^{p}\times\mathbb{R}$) feature-label pairs with some distribution function $P_{Z}^{(0)}$. At each future time $t$, we observe a test point’s features $X_{n+t}$ and are interested in rigorous UQ for an ML estimate of the unknown true label $Y_{n+t}$. For simplicity we interpret $t$ as a timestep, but in general each datapoint only needs an identifying index; also, we often let $t=1$ to focus on a single test point indexed $n+1$. When distinction is needed, capital letters (e.g., $Z_{i}$) denote random variables, and lower case (e.g., $z_{i}$) denote observed values.
2.1 Standard Conformal Prediction
Conformal prediction (CP) is a framework for predictive inference: the problem of computing a predictive confidence set $\widehat{\mathcal{C}}_{n}(x)$ (in regression, this is often an interval) likely to contain the true label at least at a user-specified target rate $1-\alpha\in(0,1)$. This desiderata is called valid coverage: [^2] We focus on marginal coverage (i.e., on average over many repeated experiments) rather than conditional coverage; see Foygel Barber et al. (2021) for more details on this distinction.
$\mathbb{P}[Y_{n+1}\in\widehat{\mathcal{C}}_{n}(X_{n+1})]\geq 1-\alpha.$ (1)
For an accessible introduction to CP see Angelopoulos et al. (2022) or for the definitive reference see Vovk et al. (2005). Computing conformal predictive sets requires access to a real-valued score function $\mathcal{S}$ that takes as its inputs a point $(x,y)$ and a “bag” [^3] That is, a multiset. CP methods assume the score-fitting algorithm is symmetric with respect to permutations of its training inputs $\bar{Z}$; we maintain this assumption unless stated otherwise. of other examples $\bar{Z}$, where $\mathcal{S}\big{(}(x,y),\bar{Z}\big{)}$ can be thought of as quantifying how “nonconforming” or “strange” the point $(x,y)$ is relative to $\bar{Z}$. For example, the absolute value residual $|\widehat{\mu}_{\bar{Z}}(x)-y|$ is a common score function, where $\widehat{\mu}_{\bar{Z}}$ is an ML predictor trained on the examples $\bar{Z}$. For shorthand, we will sometimes write the fitted score function as $\widehat{\mathcal{S}}(x,y)=\mathcal{S}\big{(}(x,y),\bar{Z}\big{)}$ and outputted score values as $V_{i}=\widehat{\mathcal{S}}(Z_{i})$. A standard CP set $\widehat{\mathcal{C}}_{n}(x)$ is then constructed by taking a (conservative) quantile on the empirical distribution of score values for each candidate $y\in\mathcal{Y}$
$\widehat{\mathcal{C}}_{n}(x)=\Big{\{}y\in\mathcal{Y}:\widehat{\mathcal{S}}(x,y)\leq\mathrm{Q}_{1-\alpha}\big{(}V_{1:n}\cup\{\infty\}\big{)}\Big{\}},$ (2)
where $Q_{1-\alpha}$ is an empirical quantile evaluated at level $1-\alpha$.
Standard CP methods assume that all the data $Z_{1},\dots,Z_{n+1}$ are exchangeable—that is, that their joint distribution is invariant to permutations, a special case being independent and identically distributed (IID). Under exchangeability, $\widehat{\mathcal{C}}_{n}(x)$ in Eq. (2) achieves the valid coverage guarantee in Eq. (1). This standard CP coverage guarantee is distribution-free in the sense of making no assumptions about the specific parametric form of the joint data distribution—this is not to be confused with robustness to distribution shift, which violates exchangeability for example when the distribution of $Z_{n+1}$ differs from that of $Z_{1:n}$. Exchangeability is thus still a strong assumption in its own way. For instance, it excludes many environment-dependent or time-dependent distribution shifts which often occur in practical settings.
2.2 Weighted Conformal Prediction for Covariate Shift
Tibshirani et al. (2019) presented an alternate proof of the conformal coverage guarantee in a way that isolated the role of exchangeability and allowed them to generalize conformal prediction to a broader class of distributions they termed weighted exchangeable. In particular, they thoroughly developed a specific weighted CP algorithm for the standard covariate shift setting (Shimodaira, 2000), where the marginal distribution of the inputs $X$ may differ independently between training and test distributions, but the conditional distribution $Y\mid X$ is assumed to be invariant:
[…]
$(X_{n+1},Y_{n+1})$ $\hskip 1.42262pt\sim P_{Z}^{(1)}=P_{X}^{(1)}\times P_{Y\mid X},\ \text{independently}.$
Note that in standard covariate shift the test distribution $P_{Z}^{(1)}$ is independent from the training data; it is fixed a priori and cannot change depending on different draws of $Z_{1:n}$. Tibshirani et al. (2019)’s weighted CP algorithm for standard covariate shift generalizes Eq. (2) with likelihood ratio function weights $w(x)=dP_{X}^{(1)}(x)/dP_{X}^{(0)}(x)$.
Fannjiang et al. (2022) followed the general approach of Tibshirani et al. (2019) to derive a specific weighted CP algorithm with a valid coverage guarantee (Eq. (1)) in a setting they call (one-step) feedback covariate shift (FCS), where $P_{X;Z_{1:n}}^{(1)}$ is explicitly dependent on the realized draw $Z_{1:n}$. While this work was an important extension of CP to dependent data, they did not consider a more general “multistep” FCS case that could describe how data shifts may accumulate over time, for instance in active learning. Their focus on the one-step FCS case could be because they posit as the basis of their theoretical analysis an assumption they call pseudo-exchangeability, which explicitly characterizes one-step FCS but does not formally describe its multistep analog (see Appendix C.3), which we introduce next.
2.3 Multistep Feedback Covariate Shift
Whereas standard and feedback covariate shift are “one-step” distribution shifts between training and test data batches, in this work we focus our practical implementations on a more general setting we call multistep feedback covariate shift (MFCS). MFCS is not only a clear example of a sequentially-dependent data distribution that goes far beyond exchangeability, but it moreover characterizes important instances of feedback-loop shifts inherent to the actions of AI/ML agents. For each time $t\in\{1,...,T\}$, MFCS allows $X_{n+t}$ to change arbitrarily depending on the past data $Z_{1:(n+t-1)}=\{Z_{i}\}_{i=1}^{n+t-1}$, while assuming that $Y\mid X$ remains invariant. More precisely, MFCS assumes
[…]
$(X_{n+t},Y_{n+t})$ $\hskip 1.42262pt\sim P_{X;Z_{1:(n+t-1)}}^{(t)}\times P_{Y\mid X},\;t=1,\dots,T.$
Note that $n$ is the initial sample size of an IID (or exchangeable, more generally) initialization, while $t$ is the number of non-exchangeable MFCS steps past this initialization. Setting $T=1$ immediately reduces Eq. (3) to one-step FCS as considered by Fannjiang et al. (2022) and moreover to standard covariate shift if independence is further assumed.
MFCS is commonly induced by ML agents whenever the ML system is able to actively select the next datapoint $X_{n+t}$ (e.g., based on a cost or utility criterion), informed by its prior experience $Z_{1:(n+t-1)}$, so long as the predictive goal or “concept” of interest (i.e., $Y\mid X$) remains the same. For instance, MFCS describes active learning, where the goal is to efficiently construct a training corpus to improve model performance, as well as black-box optimization, where the ML model is an instrument to optimize $X$ with respect to some unknown cost (or utility) function that is only indirectly accessible by each observation of $Y$.
3 Related Work
Conformal prediction under distribution shift. Prior work can broadly be categorized into variants of weighted conformal prediction (Tibshirani et al., 2019; Podkopaev & Ramdas, 2021; Fannjiang et al., 2022; Prinster et al., 2022, 2023; Barber et al., 2023; Farinhas et al., 2023) and adaptive conformal prediction (Gibbs & Candes, 2021; Zaffran et al., 2022; Angelopoulos et al., 2023; Feldman et al., 2023; Bhatnagar et al., 2023). The former approach exploits knowledge or estimates of the distribution shift to proactively adjust inference and thereby achieve guarantees that hold at inference time; meanwhile, the latter aims to maintain a target risk threshold by retroactively responding to threshold violations and thereby obtain guarantees most meaningful “in the long run” rather than at a given inference time.
This work builds on the literature of weighted CP, and our practical applications focus on cases of MFCS where the shift is agent-induced and thus known. Aside from Tibshirani et al. (2019) and Fannjiang et al. (2022), Prinster et al. (2022) and Prinster et al. (2023) also studied CP under (one-step) standard or feedback covariate shift by extending cross-validation-style CP methods (from Barber et al. (2021)) to those settings for flexibly balancing compute-versus-sample efficiency. Otherwise, our work is most closely related to Barber et al. (2023), which bounds the worst-case “coverage gap” (i.e., loss of coverage below Eq. 1) if the conformal weights are unknown and instead fixed a priori independently of the data. In contrast, our extensions of CP to non-exchangeable distributions use data-dependent weights to achieve coverage validity guarantees as in Eq. 1 (without a gap in coverage).
Conformal inference for decision-making. There is a growing body of work applying the ideas of CP to decision-making to inform utility estimates (Vovk & Bendtsen, 2018; Stanton et al., 2023; Salinas et al., 2023) or satisfy constraints (Zhang et al., 2023; Dixit et al., 2023; Laufer-Goldshtein et al., 2023; Jin & Candès, 2023). In particular, Stanton et al. (2023) drew motivation from Fannjiang et al. (2022) to combine CP with Bayesian optimization but noted that a gap in theory prevented a formal coverage guarantee after the first round, a gap we address in this work.
4 Theoretical Contributions and Methods
4.1 The Role of (Weighted) Exchangeability and Related Assumptions in Conformal Prediction
Informal Statement of the Key Insight One can think of conformal prediction (CP) as an inverted permutation test (with a conservative adjustment). Informally, our first key insight is that—from this perspective of an inverted permutation test—the only role of exchangeability, weighted exchangeability (Tibshirani et al., 2019), or related distributional assumptions such as pseudo-exchangeability (Fannjiang et al., 2022) in CP is to enable tractability by telling us how much to “weigh” or “count” a given permutation of datapoints (or score values); in particular, by facilitating computation of the relative likelihood of each sequence of observations, relative to some core function (that does not depend on the ordering if its inputs), given the unordered “bag” of values $\{z_{1},...,z_{n+1}\}$ (see Appendix C details on these definitions and their roles). So, extensions of CP should be possible so long as they are able to quantify the relative likelihood of any ordering of datapoints (or scores).
Formal Sketch of the Key Insight We now sketch with formal notation our main insight. This sketch will correspond to viewing CP as an inverted permutation test, provide intuition for our main theorem in Section 4.2, and moreover highlight key ideas underlying the theorem’s first proof step.
Constructing CP from an inverted permutation test begins by conditioning on the set [^4] For simplicity, we assume all the values are distinct to work with sets rather than multisets; in general, the argument holds with more complex notation using uniform randomness to break ties. of data values $\{z_{1},...,z_{n+1}\}$—that is, let $E_{z}$ denote the event that $\{Z_{1},...,Z_{n+1}\}=\{z_{1},...,z_{n+1}\}$. To be clear, this notation means that the unordered “bag” of values is known but not whether $Z_{i}=z_{i}$, and so on; that is, we know the random variable $Z_{i}$ takes on an observed value in $\{z_{1},...,z_{n+1}\}$, but we do not yet know which value. Conditioning on $E_{z}$, we are then interested in the probability that the test point’s score $V_{n+1}$ was actually $v_{i}$, for each $i\in\{1,...,n+1\}$, which we can write as
$\mathbb{P}\{V_{n+1}=v_{i}\ |$ $E_{z}\}=\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$
$=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+1)})},$ (4)
where $f$ denotes the joint probability density function (PDF). Standard CP proceeds directly from this statement by leveraging exchangeability to show that every permutation of the data values is equally likely, thus reducing the problem to counting permutations; weighted CP proceeds by factorizing $f$ using weighted exchangeability, with the product of weight functions $\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})$ telling us how much weight to give to the permutation $\sigma$; and, CP with pseudo-exchangeable data proceeds by factorizing $f$ similarly, with more explicit weight functions describing a specific instance of potential dependencies (See Appendix C.3).
More generally, the key insight comes from observing that equation (4) does not rely on any assumptions about the joint distribution $f$. This can be seen by deriving the statement using Bayes rule and the law of total probability (LOTP):
$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ =$=\frac{\mathbb{P}\{Z_{n+1}=z_{i},E_{z}\}}{\mathbb{P}\{E_{z}\}}$
$=\frac{\sum_{\sigma:\sigma(n+1)=i}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}.$
That is, the last step uses the LOTP in both the numerator and the denominator—for instance, in the denominator the summation is over all the possible ways that the event $E_{z}$ could occur, which amounts to all the different possible permutations $\sigma$ of the values. The crux of extending CP to any, possibly non-exchangeable joint distribution can thus reduce to the problem of either computing $f(z_{\sigma(1)},...,z_{\sigma(n+1)})$ directly for any ordering $\sigma$ of the observations, or to be able to factorize $f$ into terms that are tractable to compute (e.g., weight or factor functions) and remaining terms that do not depend on $\sigma$ (e.g., a core measure that is invariant to permutations and thus cancels out in the ratio).
4.2 A General View of Conformal Prediction
Main Formal Result With the insight from Section 4.1 in mind, we are now ready to see how—in theory—extensions of CP are possible for any, potentially non-exchangeable distribution [^5] To be precise, we require a non-singular distribution so that there exists a valid PDF $f$, or more generally, a valid Radon-Nikodym derivative with respect to an arbitrary base measure. of the calibration and test data $Z_{1},...,Z_{n+1}$ (we let $t=1$ for simplicity, but in general our argument holds for $Z_{1},...,Z_{n+t}$). This section’s theorem formally states this result as the coverage guarantee for a generalized weighted CP set for any joint probability density function (PDF) $f$. We will then soon discuss both the practical benefits and limitations of such a general view.
Let us define nonconformity scores more explicitly as
$\begin{cases}V_{i}^{(x,y)}&=\mathcal{S}(Z_{i},Z_{1:n}\cup\{(x,y)\}),\ i\in\{1,...,n\},\\ V_{n+1}^{(x,y)}&=\mathcal{S}\big{(}(x,y),Z_{1:n}\cup\{(x,y)\}\big{)},\end{cases}$ (5)
and for condensed notation, let us denote $\mathbb{P}_{n+1}\{z_{i}|E_{z}\}=\mathbb{P}\{Z_{n+1}=z_{i}|E_{z}\}$, recalling Eq. (4). To enable us to write a weighted empirical distribution of values with probability weights $\mathbb{P}_{n+1}\{z_{i}|E_{z}\}$ applied to each value $v_{i}$, we will also now denote use $\delta_{v}$ to denote a point mass at the value $v$ (e.g., in the special case of uniform weights $\tfrac{1}{m}$ over values $v_{1},...,v_{m}$, then $Q_{1-\alpha}(v_{1:m})=Q_{1-\alpha}(\sum_{i=1}^{m}\tfrac{1}{m}\delta_{v_{i}})$). We now have the notation needed for our main theorem.
Theorem 4.1.
Assume that $Z_{i}=(X_{i},Y_{i})\in\mathbb{R}^{d}\times\mathbb{R},i=1,...,n+1$ have the joint PDF $f$. For any measurable score function $\mathcal{S}$, and any $\alpha\in(0,1)$, define the generalized conformal prediction set (based on $n$ calibration samples) at a point $x\in\mathbb{R}^{d}$ by
$\widehat{\mathcal{C}}_{n}(x)=\bigg{\{}y\in\mathbb{R}:V_{n+1}^{(x,y)}\leq$ (6)
$Q_{1-\alpha}\bigg{(}\sum_{i=1}^{n}\mathbb{P}_{n+1}\{Z_{i}|E_{z}\}\delta_{V_{i}^{(x,y)}}+\mathbb{P}_{n+1}\{Z_{n+1}|E_{z}\}\delta_{\infty}\bigg{)}\bigg{\}}$
where $V_{i}^{(x,y)},i\in\{1,...,n+1\}$ are as in (5) and $\mathbb{P}_{n+1}\{z_{i}|E_{z}\}$ is notation for (4). Then, $\widehat{\mathcal{C}}_{n}$ satisfies
$\mathbb{P}\big{\{}Y_{n+1}\in\widehat{\mathcal{C}}_{n}(X_{n+1})\big{\}}\geq 1-\alpha.$
Proof Sketch We defer the full proof for the result to Appendix A.1, but a sketch of the proof has two main steps:
  1. 1.
    CP as an inverted (weighted) permutation test: With setup as in Section 4.1, use Eq. (4) to derive a general quantile lemma, which informally states the following: For $\beta\in(0,1)$, the test point’s score $V_{n+1}$ is contained within the level $\beta$ quantile of the weighted empirical distribution of score values (with weights defined as in Eq. (4)) with probability at least $\beta$.
  2. 2.
    Conservative adjustment to connect to general CP set: Connect the general quantile lemma from the first proof step to general CP set (Eq. (6)) by conservatively adjusting for the fact that the true test point’s score value is unknown. That is, replacing the test point’s score value with $\infty$ in the empirical distribution maintains validity, and setting $\beta=1-\alpha$ gives the theorem.
Remark 4.2.
The score functions in Eq. (5) correspond to the “ordinary” full CP method, but the same proof argument and result follows for the “deleted” full CP method (Vovk et al., 2005), which uses leave-one-out scores to prevent overfitting. Moreover, the same result follows for split conformal prediction as a special case. (See Appendix A.2.)
Remark 4.3.
As written, Theorem 4.1 maintains the convention of requiring $\mathcal{S}$ to be symmetric with respect to any training data also used in calibration. This condition is relevant for full CP methods, but for split CP, this symmetry holds trivially due to training and calibration sets being separate. (See Appendix A.3 for further details and discussion.)
Limitations and Benefits of the General View of CP There are two main obstacles to computing Eq. (6) in practice: a computational complexity challenge and an epistemic challenge regarding knowledge about $f$. That is, in the general case for an arbitrary or unfactorized $f$, Eq. (4) requires $\mathcal{O}((n+1)!)$ evaluations of $f$, which quickly becomes intractable. Secondly, we often do not know $f$; indeed, the primary motivation of distribution-free UQ is that very deficiency. Exchangeability, weighted exchangeability (Tibshirani et al., 2019), and related conditions such as pseudo-exchangeability (Fannjiang et al., 2022) introduce distributional assumptions which alleviate these challenges.
One immediate benefit of the general presentation of conformal prediction in Theorem 4.1 is a shift in perspective: because the theorem holds for any joint distribution $f$, for any particular $f$ the question is not whether a CP set with valid coverage exists, it is instead whether we can compute the prediction set in practice. Consequently—from this general perspective—“quasi-exchangeability” definitions are no longer necessary conditions for CP theory; rather, the only role of these definitions is to enable practical tractability by simplifying and eliminating terms from the computation (of Eq. (4))—for example, at one extreme, exchangeability “assumes away” $f$ from the computation entirely. (Appendix C provides further formal details on this point, including how previous CP validity guarantees can be viewed as corollaries to Theorem 4.1.) Specific weighted CP algorithms with corresponding coverage validity guarantees can thus be derived for any (possibly non-exchangeable) distribution. In section 4.3, we outline how to perform such derivations in general and provide the MFCS case as a worked example.
Further Discussion of Tibshirani et al. (2019) This general view of CP is implicit in a close reading of Tibshirani et al. (2019), however the authors did not explicitly state or prove the general result we present in Theorem 4.1. Instead, Tibshirani et al. (2019) premised their main result on their definition of weighted exchangeability despite this condition not being strictly necessary for CP in the broad sense. While we maintain that weighted exchangeability is still a useful definition (see Appendix C.2), subsequent literature has seemingly been led to believe that new definitions of “quasi-exchangeability” are necessary for advancing and delineating the boundaries of CP theory—for example, Fannjiang et al. (2022)’s introduction of pseudo-exchangeability to study the single-step FCS setting. We hope our contribution will lead to wider recognition of the generality of Tibshirani et al. (2019)’s analysis and foster a broadened scope of possibility in the study of CP: in particular, we encourage a shift away from viewing “quasi-exchangeability” conditions as theoretical boundaries to be overcome, and instead towards primarily using such conditions to clearly specify the assumptions granted by a given setting of interest or adopted for practical tractability.
4.3 Deriving Weighted CP Algorithms for MFCS, or Any Data Distribution
A General Procedure for Deriving Weighted CP Algorithms with Valid Coverage Guarantees To improve the accessibility and practical utility of our main theoretical contribution, we now outline a general procedure that can be used to derive weighted CP algorithms for any specific joint PDF $f$ over the calibration and test data. Theorem 4.1 implies the existence of this procedure without premising it on any assumptions on $f$, while emphasizing that the only role of such assumptions is for tractability in the last step.
  1. 1.
    List assumptions (if any): For a given application setting, specify assumptions on $f$, if any, such as conditional independence or invariance assumptions. (For example, the MFCS assumptions are formally given by Eq. (3); informally, MFCS assumes $Y|X$ is invariant, while $X$ is dynamic depending on past observations.) Alternatively, one could begin by describing an application’s data-generating process by a probablistic graphical model (e.g., a Markov random field, Bayesian network, or a causal directed acyclic graph) and use the conditional independence assumptions implied by that graphical model (for introductions to such models see Koller & Friedman (2009); Pearl (2009)). [^6] Probablistic graphical models can intuitively represent a data-generating process as a graph with nodes corresponding to random variables (or random vectors). Missing edges in such a graph imply conditional independences in the joint PDF $f$; the converse depends on further assumptions (Pearl, 2009).
  2. 2.
    Factorize $f$: Factorize the joint PDF $f$ using standard probability rules (e.g., Bayes or chain rule) and assumptions from Step 1 to separate dynamic factors from those that are invariant (to permutations $\sigma$ of the data indices). (For example, for MFCS $f$ can be written as a product of dynamic factors $X_{j}\mid Z_{1},…,Z_{j-1}$ and invariant factors $Y\mid X$.)
  3. 3.
    Compute or estimate weights: Plug the factorized form of $f$ from Step 2 into Eq. (4) to obtain the normalized calibration and test point weights, and simplify by canceling out permutation-invariant and constant factors. Compute or estimate the simplified weights, and plug them into (6) to obtain the CP set. If computation is performed exactly, then the resulting CP set has a valid coverage guarantee premised on the assumptions from Step 1 (as a corollary to Theorem 4.1).
Deriving CP for Agent-Induced MFCS For a concrete demonstration of how this general procedure can be used to derive novel CP methods, we now return to the MFCS setting introduced in Section 2.3, which characterizes common instances of agent-induced feedback-loop data shifts. We provide only a sketch of key ideas here; for full details see Appendix B. Step 1 of this procedure for MFCS begins by recalling Eq. (3), the conditional independence or invariance assumptions that we used to define MFCS. Using standard probability rules and leveraging these MFCS assumptions, for Step 2 we are able to factorize the joint PDF $f$ to separate the time-dependent factors corresponding to $X_{j}\mid Z_{1},…,Z_{j-1}$ from the time-invariant factors corresponding to $Y\mid X$:
$f($ $z_{1},...,z_{n+t})=$
$\prod_{j=1}^{n+t}\Big{[}\underbrace{\mathbb{P}(x_{j}\ |\ z_{1},...,z_{j-1})}_{\text{Time-dependent factors}}\Big{]}\cdot\underbrace{\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(y_{j}|x_{j})\Big{]}}_{\text{Time-invariant factor}}.$ (7)
We focus on time dependence and invariance here only because we assume that the data indices indicate timesteps; in general, Step 2 aims to separate terms that depend on permutations of the data indices from those that are invariant to such permutations (if any). Lastly, for Step 3, we use our factorization of $f$ to simplify the CP calibration and test point weights. To do so, we slightly modify our prior notation by letting $E_{z}^{(t)}$ denote the event $\{Z_{1},...,Z_{n+t}\}=\{z_{1},...,z_{n+t}\}$; then, plugging Eq. (7) into the analog of Eq. (4) for test point $n+t$ yields
[…]
$=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\mathbb{P}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}\mathbb{P}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})},$ (8)
after canceling out $\prod_{j=1}^{n+t}\big{[}\mathbb{P}(y_{\sigma(j)}|x_{\sigma(j)})\big{]}$ from the numerator and denominator, as it is invariant to permutations $\sigma$.
As a corollary to Theorem 4.1, a CP method with weights given by Eq. (4.3) has a valid coverage guarantee premised on the MFCS assumptions in Eq. (3). We can thus turn to practical estimation. Firstly, in the agent-induced MFCS settings that we focus on in our experimental evaluations (Section 5), the epistemic challenge is overcome because the dynamic components of $f$ are known entirely. That is, the invariant factor corresponding to $Y|X$ cancels in the ratio, and in our practical evaluations the remaining factors $\mathbb{P}(x|Z_{1},...,Z_{j-1})$ represent ML-agent-controlled query probability functions at each timestep $j$. Next, we examine computational complexity. Due to the MFCS IID initialization implying $\mathbb{P}(x|Z_{1},...,Z_{j-1})=\mathbb{P}(x)$ for $j\leq n$, and when $\mathbb{P}(x|Z_{1},...,Z_{j-1})$ is computed from an ML model that treats $Z_{1},...,Z_{j-1}$ symmetrically (as in our experiments), the exact computation of Eq. (4.3) has complexity $\mathcal{O}(\prod_{j=1}^{t}(n+j))$. Though reduced from $\mathcal{O}((n+t)!)$, the complexity for arbitrary $f$, this this still quickly becomes intractable for large $t$. To alleviate this remaining complexity bottleneck, we propose estimating Eq. (4.3) by using only the “highest-order” terms from the $d$ most recent timesteps. That is, for a user-specified estimation depth $d\in[t]$, then we define our “$d$-step” estimate of Eq. (4.3) as
$\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}=\widehat{p}^{\ (d)}\{Z_{n+t}=z_{i}\mid E_{z}^{(t)}\}$ (9)
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{j=\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n+t+1-d}\mathbb{P}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod^{n+t}_{j=\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n+t+1-d}\mathbb{P}(x_{\sigma(j)}\mid z_{\sigma(1)},...,z_{\sigma(j-1)})},$
which reduces the complexity to $\mathcal{O}(\prod_{j=t+1-d}^{t}(n+j))$, a tractable polynomial when $d$ is small. Appendix B.2 provides a derivation for our specific MFCS CP algorithms, which use a recursive implementation for further efficiency.
Figure 1: Results for fluorescent protein design ($n=32$, $\lambda=8$) with a linear ridge regressor, comparing the proposed MFCS full CP methods to standard full CP and one-step weighted full CP baselines. We repeated this experiment with 1,000 random seeds; the error bars for mean coverage and mean predicted fitness represent standard error, while the error bars for median interval width represent interquartile range. The proposed MFCS full CP maintains coverage where the baselines do not. Its intervals are wider than those of the one-step FCS baseline where the latter loses coverage ($t=2,3,4$), but similar where that baseline maintains coverage ($t=1,5$).
5 Experimental Results
We evaluate the practical performance of our proposed CP methods for ML-agent-induced MFCS on synthetic black-box optimization (specifically protein design) and active learning tasks. All experiments focus on a regression setting with all CP methods using the absolute residual score $\widehat{\mathcal{S}}(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}_{t}(x)$ is the predicted mean. Our proposed methods are defined by estimating Eq. (6) with weights as in Eq. (9). The primary desired criterion for evaluating a method’s reliability is whether its prediction intervals maintain empirical coverage at or above the target coverage level $1-\alpha$. Secondarily—if target coverage is achieved—then smaller predictive sets are more informative (i.e., predicting the full label space $\widehat{\mathcal{C}}_{n}(x)=\mathcal{Y}$ could trigger deferral to an expert but does not itself convey new insights from the predictor). In all experiments we compare to Exchangeable CP (Vovk et al., 2005) and One-Step FCS CP (Fannjiang et al., 2022), and in the active learning experiments we additionally compare with Adaptive Conformal Inference (ACI) (Gibbs & Candes, 2021). Our proposal distributions follow the form $\mathbb{P}_{t}(X_{n+t}=x):=\mathbb{P}(X_{n+t}=x\mid Z_{1:(n+t-1)})\propto\exp(\lambda\cdot u_{t}(x))$, where $u_{t}$ is some utility function and $\lambda\geq 0$ is a hyperparameter that controls the magnitude of agent-induced covariate shift. In all experiments we corrupt observed labels with Gaussian noise to simulate noisy observations. We add every query point to the training set for the full CP experiments, and for split CP we randomly assign the query points to either training or calibration by flipping a fair coin. (Details in Appendix D.)
5.1 Multistep Protein Design Experiments
Optimization-Induced MFCS Experimental Details Our protein design procedure is similar to a common approach in the biomolecular engineering literature (Biswas et al., 2021; Zhu et al., 2024) and related works in CP (Fannjiang et al., 2022; Prinster et al., 2023), but extended to a more realistic online setting where protein sequence proposals and model updates are interleaved and dependent (Stanton et al., 2022; Gruver et al., 2023; Angermueller et al., 2023). All methods start with an IID initialization by sampling $n$ points uniformly at random from a combinatorially complete library (Poelwijk et al., 2019), splitting the data equally into training and calibration sets if using split CP. See Appendix D.1 for additional details. We evaluate the mean coverage and median intervals width as a function of the number of optimization steps past initialization, fitting the regressor to $Z_{1:n+t-1}$, and sampling one query point with replacement (with added measurement noise) from the library with probability $\mathbb{P}_{t}(X_{n+t}=x)\propto\exp(\lambda\cdot\widehat{\mu}_{t}(x))$.
Figure 2: Results for fluorescent protein design with an MLP regressor ($n=64$, $\lambda=5$), comparing the proposed MFCS Split CP methods to standard split CP and one-step weighted split CP baselines. The experiment is repeated with 500 random seeds; the error bars for mean coverage and mean predicted fitness represent standard error, while for median interval width they represent the interquartile range. Error bars extending beyond the top of the figure indicate that the upper quartile is infinite. MFCS Split CP maintains coverage, but its interval are occasionally very wide, suggesting the proposal distribution is too aggressive for consistently informative uncertainty quantification.
Full CP Results for Multistep Protein Design Our first evaluations focus on variants of Full CP (with leave-one-out scores; see Appendix A.2), which efficiently uses the same dataset for both training and calibration (often resulting in more accurate and sharper prediction intervals relative to other CP methods). Full CP’s data efficiency comes at a steep computational cost, with its complexity corresponding to the number of predictors that must be trained: For an arbitrary $\widehat{\mu}$, computing the MFCS Full CP weights in Eq. (4.3) requires training $|\mathcal{Y}|\cdot\prod_{j=1}^{t}(n+j)$ distinct predictors, where $|\mathcal{Y}|$ is the cardinality of the label space. [^7] In practice, regression labels are discretized by a fine grid. If $\widehat{\mu}$ is a ridge regression predictor, for the training cost is reduced to $\prod_{j=1}^{t}(n+j)$, and we moreover set our estimation depth to $d=2$ for computing our estimated MFCS weights Eq. (9), which reduces the training cost to $(n+t)\cdot(n+t-1)$.
Figure 1 reports the results for our proposed MFCS Full CP method (blue circles) with hyperparameters reported in the caption. On both the blue and red fluorescent protein design datasets, and on all five evaluated design steps, our proposed MFCS Full CP method maintains coverage above the target level even when both baselines lose coverage below this level. The MFCS method’s coverage is maintained by increasing the conservativeness (size) of its prediction intervals relative to the One-Step baseline after the first design step. The superimposed fitness curves on the right verify that differences in coverage can be attributed to the different CP methods, not the data being collected.
Figure 3: Active learning experiments of proposed multistep split CP method for $d=3$ (red) compared to baselines of unweighted split CP (orange), one-step split CP (green), and ACI (gray). The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. For all experiments, a holdout test set of 250 samples was first sampled uniformly at random from the full dataset to track the accuracy of the base predictor as the active learning progressed. To simulate a practical active learning setting where there often tends to be sample-selection bias in how the initial training data is obtained relative to the desired test distribution, we sampled the initial training and calibration datasets from the full dataset with probabilities proportional to $\exp(\gamma\cdot X_{\text{PCA}1}^{\text{normed}})$, where $X_{\text{PCA}1}^{\text{normed}}$ is the min-max-normed first principal component representation of the data. All values are computed over 350 distinct random seeds.
Efficient Computation with Split CP for Multistep Protein Design While the high training cost of Full CP is prohibitive for many types of ML predictors, this limitation is avoided by variants of Split CP, which we also extend to MFCS. Variants of Split CP reduce computational effort by splitting the available data into a “proper” training set that is only used to fit a single ML model, and a separate holdout “calibration” set that is evaluated with the fitted model (via the score function) to construct the CP set. Split CP’s computational complexity thus corresponds to the evaluation cost of the pre-fitted model on the calibration and test data.
The results in Figure 2 use a neural network regressor on the protein design datasets to evaluate our MFCS Split CP method. We moreover compare the effect of several estimation depths $d$ for estimating the MFCS weights as in Eq. (9). For weight estimation depths $d\in\{3,4\}$, our proposed MFCS CP methods maintain target coverage across all 10 evaluated design steps, for both the blue and red fluorescent protein design datasets, even when the baselines break down. Increasing the estimation depth tends to improve coverage at later design steps, for instance relative to the one-step FCS, apparently with diminishing returns (the red 3-step and purple 4-step methods have nearly overlapping performance). Coverage is maintained by the MFCS Split CP methods by slightly increasing the size of the intervals relative to the One-Step baseline at later design steps. However, both the MFCS Split CP method and the One-step FCS Split CP method occasionally have noninformative (infinite width) intervals, which could be unacceptable in some application settings. In Appendix E, we show that adaptively bounding the ML agent’s query probability function can achieve consistently informative (finite) interval widths without losing coverage, by restricting the ML agent’s feedback shift.
5.2 Active Learning Induced MFCS
Now we turn to an active learning setting where MFCS is induced by a max-entropy data acquisition strategy, where the ML agent selects new queries for annotation by maximizing its (Bayesian) predictive uncertainty. In this experiment we use a Gaussian process (GP) regressor as the predictor and take $u_{t}(x)=\mathbb{V}[f(x)|Z_{1:n+t-1}]$ (i.e. the GP posterior variance). See Appendix D.2 for model and dataset details.
Active Learning Experimental Results Figure 3 demonstrates the performance of our proposed MFCS Split Conformal methods for several depths of weight approximation. Whereas the baseline methods lose coverage, the proposed MFCS Split CP methods maintain target coverage for all datasets and all of the 70 evaluated active learning steps. As the active learning process reaches later iterations where more labeled data is available for both training and calibration, the proposed MFCS Split CP methods are still able to maintain target coverage with sharper intervals.
6 Discussion
In this work we have demonstrated that weighted conformal prediction can theoretically be generalized to any joint data distribution. Exchangeability assumptions are only necessary to rigorously specify the conditions under which a specific practical algorithm (e.g. standard CP) is equivalent to the more general formulation, which is always valid but may not be computable. Promising avenues for future work include extending our general marginal coverage guarantees to more general loss functions (Angelopoulos et al., 2022) and identifying a minimal set of assumptions for good conditional coverage in practice (Gibbs et al., 2023).
We have also provided a practical algorithm for CP in the presence of multiple rounds of agent-induced, data-dependent covariate shift by straightforward derivation, allowing us to maintain coverage on data collected by active learning and black-box optimization agents. Finally we have shown that these agents can attain both empirically valid coverage and finite interval widths by restricting the degree of shift the agent can introduce. These findings bring us to the control/consistency dilemma at the heart of statistical decision-making. To control is to change, but statistics requires consistency. Like the explore/exploit dilemma, the right tradeoff depends on the situation at hand. As machine learning agents assume increasing levels of responsibility in gravely consequential settings, it is more important than ever that we face the control/consistency dilemma head-on.
Code Availability
GitHub repository with code for all experiments: https://github.com/drewprinster/conformal-mfcs
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work which we do not think must be specifically highlighted here. However, we note that our contributions are specifically aimed at improving the reliability of methods for uncertainty quantification of machine learning predictions, which is often critical for managing risks and improving the trustworthiness of such systems.
Acknowledgements
D.P. was partially funded by a fellowship from the Johns Hopkins University (JHU) Mathematical Institute for Data Science (MINDS); D.P. and S. Saria were partially funded by National Science Foundation (NSF) grant #1840088; D.P., A.L., and S. Saria were partially funded by the Gordon and Betty Moore Foundation grant #12128. A.L. was also partially funded by a JHU Discovery Award and a JHU Institute for Assured Autonomy (IAA) seed grant. The authors would like to thank Anastasios Angelopoulos for early discussions that later helped inform this work and Clara Fannjiang for helpful feedback on a manuscript draft.
Appendix A Main Result Proof and Details
A.1 Proof of Theorem 4.1
In this section we present the proof for our main result given in Theorem (4.1). Our proof builds on ideas in Tibshirani et al. (2019)’s alternate proof of their Lemma 1 (for conformal prediction assuming exchangeable data) and their proof of their Lemma 3 (for conformal prediction assuming weighted exchangeable data). However, those proofs make key substitutions using the definitions of exchangeability and weighted exchangeability respectively, which thereby yields guarantees that are premised on those definitions. In contrast, the proof we present here makes no such substitution, and thus holds for an arbitrary joint probability density function $f$. We will call attention to the key step where our proof critically differs from those in Tibshirani et al. (2019) in substance, and even for steps that we have in common with Tibshirani et al. (2019) we aim to provide more explicit derivation explanations for increased accessibility (e.g., where explicit step justifications may be omitted in Tibshirani et al. (2019)).
As we sketch in Section 4.1 of our main text, our proof corresponds to a view of CP as an inverted permutation test. Let $E_{z}$ denote the event of observing the unordered set [^8] As in Tibshirani et al. (2019) we assume for simplicity that the score values $V_{1},...,V_{n+1}$ are distinct almost surely, though the result also holds in more general cases (but with more complicated notation involving randomness from uniform random variables). of our data values. In particular, $E_{z}$ denotes the event that $\{Z_{1},...,Z_{n+1}\}=\{z_{1},...,z_{n+1}\}$, but importantly, the lack of ordering means that this does not imply that $Z_{i}=z_{i}$ for all $i\in\{1,...,n+1\}$; we do not yet know whether $z_{1}$ is the value obtained by $Z_{1}$, or by $Z_{2}$, or by $Z_{n+1}$, and so on. Similarly, let $E_{v}$ denote the event $\{V_{1},...,V_{n+1}\}=\{v_{1},...,v_{n+1}\}$ for the scores. Also, recall the simplifying notation $v_{i}=\mathcal{\widehat{S}}(z_{i})$ for the value computed by our fitted score function $\mathcal{\widehat{S}}$ with input $z_{i}$ (where $\mathcal{\widehat{S}}(z_{i})=\mathcal{S}(z_{i},z_{1:(n+1)})$, that is $\mathcal{\widehat{S}}$ is fit on the set $z_{1:(n+1)}=\{z_{1},...,z_{n+1}\}$, the second input to $\mathcal{S}$).
With this notation, we consider the probability that, conditioned on $E_{v}$, the value corresponding to the random variable $V_{n+1}$ was actually $v_{i}$. That is, we consider
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}=\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\},\ i\in\{1,...,n+1\},$
which follows since, conditioned on $E_{z}$, there are an equal number of scores $v_{1},...,v_{n+1}$ as data points $z_{1},...,z_{n+1}$, thus implying that the fitted (symmetric) score function $\mathcal{\widehat{S}}$ induces a bijection between the discrete set of distinct data observations and the discrete set of distinct scores. [^9] To be thorough and precise, the equation $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ relies on the convention in conformal prediction that $\mathcal{S}$ is a symmetric score function (see Remark 4.3). We elaborate on this point in Appendix A.3. Then, using Bayes rule and the law of total probability (LOTP), it follows that
$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ =$=\frac{\mathbb{P}\{Z_{n+1}=z_{i},E_{z}\}}{\mathbb{P}\{E_{z}\}}$
$=\frac{\sum_{\sigma:\sigma(n+1)=i}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}$ (10)
where the last step follows by applying the LOTP in both the numerator and the denominator. For example, in the denominator the summation is over all possible permutations $\sigma$ of the value, i.e., all the different orderings in which the event $E_{z}$ could have occurred. Even more explicit notation could be to write out $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})$, to emphasize that the permutation is on the values and not on the random variables in the joint distribution $f$.
Here is where our proof differs critically from the analogous proofs in Tibshirani et al. (2019). At this point, their proof corresponding to exchangeable conformal prediction uses exchangeability to substitute in $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=f(z_{1},...,z_{n+1})$, since under exchangeability $f$ is invariant to permutations $\sigma$; meanwhile, their proof corresponding to weighted conformal prediction substitutes in $f(z_{\sigma(1)},...,z_{\sigma(n+1)})=\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{1},...,z_{n+1})$, since under their definition of weighted exchangeability $f$ factorizes into weight functions $w_{j}$ and a core function $g$ that is invariant to $\sigma$. Here, unlike in Tibshirani et al. (2019), we instead maintain the generality of Eq. (10) to prove our more general conformal prediction result that holds for any $f$.
For simpler notation, let $\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}=\mathbb{P}_{n+1}\{z_{i}\ |\ E_{z}\}$. Then, statement (10) implies that, conditioned on $E_{z}$, the random variable $V_{n+1}$ is distributed according to the weighted empirical distribution of the values $\{v_{1},...,v_{n+1}\}$, with weight $\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ at each $\delta_{v_{i}}$:
$V_{n+1}|E_{v}\sim\sum_{i=1}^{n+1}\mathbb{P}_{n+1}\{z_{i}\ |\ E_{z}\}\cdot\delta_{v_{i}}.$
Letting $Q_{\beta}$ denote a level-$\beta$ quantile function, this implies
$\mathbb{P}\bigg{\{}V_{n+1}\leq Q_{\beta}\bigg{(}\sum_{i=1}^{n+1}\mathbb{P}_{n+1}\{z_{i}\ |\ E_{z}\}\cdot\delta_{v_{i}}\bigg{)}\ \bigg{|}\ E_{z}\bigg{\}}\geq\beta.$
Due to the conditioning on $E_{z}$, this is equivalent to
$\mathbb{P}\bigg{\{}V_{n+1}\leq Q_{\beta}\bigg{(}\sum_{i=1}^{n+1}\mathbb{P}_{n+1}\{Z_{i}\ |\ E_{z}\}\cdot\delta_{V_{i}}\bigg{)}\ \bigg{|}\ E_{z}\bigg{\}}\geq\beta,$
and marginalizing yields
$\mathbb{P}\bigg{\{}V_{n+1}\leq Q_{\beta}\bigg{(}\sum_{i=1}^{n+1}\mathbb{P}_{n+1}\{Z_{i}\ |\ E_{z}\}\cdot\delta_{V_{i}}\bigg{)}\bigg{\}}\geq\beta$ (11)
We note that statement (11) can be viewed as a generalized version of both Lemma 1 and Lemma 3 in Tibshirani et al. (2019). We now connect statement (11) to the general conformal prediction set defined in Theorem 4.1 to obtain the result. To do so, we observe that
$Q_{\beta}\bigg{(}\sum_{i=1}^{n+1}\mathbb{P}_{n+1}\{Z_{i}\ |\ E_{z}\}\cdot\delta_{V_{i}}\bigg{)}\leq Q_{\beta}\bigg{(}\sum_{i=1}^{n}\Big{[}\mathbb{P}_{n+1}\{Z_{i}\ |\ E_{z}\}\cdot\delta_{V_{i}}\Big{]}+\mathbb{P}_{n+1}\{Z_{n+1}\ |\ E_{z}\}\cdot\delta_{\infty}\bigg{)},$ (12)
where the right-hand side is obtained from the left-hand side by replacing $\delta_{V_{n+1}}$ with $\delta_{\infty}$ to conservatively increase the value of the right side relative to the left. Now, with the notation $V_{i}=V_{i}^{(X_{n+1},Y_{n+1})}$ for $i\in\{1,...,n+1\}$, note that the right hand side of (12) is the same quantile as is used in the construction of the general conformal prediction interval in 4.1. That is, $Y_{n+1}\in\widehat{C}_{n}(X_{n+1})$ if and only if
$V_{n+1}\leq Q_{\beta}\bigg{(}\sum_{i=1}^{n}\Big{[}\mathbb{P}_{n+1}\{Z_{i}\ |\ E_{z}\}\cdot\delta_{V_{i}}\Big{]}+\mathbb{P}_{n+1}\{Z_{n+1}\ |\ E_{z}\}\cdot\delta_{\infty}\bigg{)}.$ (13)
Due to statement (12), the probability of the event (13) is thus lower bounded by statement (11). Letting $\beta=1-\alpha$ in the quantile function then gives the result.
A.2 Additional Details for Remark 4.2
In Theorem 4.1 and the proof in Appendix A.1 we use score functions that correspond to the “ordinary” full conformal method Vovk et al. (2005). However, the same result and proof technique would also apply to generalizing the “deleted” or “leave-one-out” full conformal method (which is computationally more expensive than ordinary full CP but tends to obtain more informative prediction sets by avoiding overfitting in its construction) Vovk et al. (2005) or the split conformal method Papadopoulos (2008). The score functions for the deleted full CP method would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{-i}\cup\{(x,y)\})$ for $i\in\{1,...,n\}$, where $Z_{-i}=Z_{1:n}\backslash Z_{i}$, and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:n}\big{)}$. For split CP, the score functions could be defined using independent training and calibration sets. Letting $Z^{0}_{1:m}=\{Z_{i}^{0}\}_{i=1}^{m}$ denote the training data and $Z_{1:n}=\{Z_{i}\}_{i=1}^{n}$ denote the calibration data, the split CP score functions would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{1:m}^{0})$ for $i\in\{1,...,n\}$ and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:m}^{0}\big{)}$.
A.3 Discussion on symmetric versus nonsymmetric score functions
In the main paper, we maintained the convention in conformal prediction of requiring that $\mathcal{S}$ be a symmetric score function (e.g., see Remark 4.3). Here, we briefly consider how our main result would need to be modified for nonsymmetric score functions, while leaving more rigorous analysis for future work. First for an example, if $\widehat{\mu}_{\bar{Z}}\leftarrow\mathcal{A}(\bar{Z})$ is an ML predictor fit on training examples $\bar{Z}$ by an algorithm $\mathcal{A}$, then the absolute value residual score $\mathcal{S}\big{(}(x,y),\bar{Z}\big{)}=|y-\widehat{\mu}_{\bar{Z}}(x)|$ is symmetric if the algorithm $\mathcal{A}$ treats its data symmetrically (i.e., if it is invariant to shuffling the ordering of its inputs); meanwhile, this score function is nonsymmetric if the fitted ML preditor output by the algorithm $\widehat{\mu}_{\bar{Z}}\leftarrow\mathcal{A}(\bar{Z})$ depends on the ordering of the elements in $\bar{Z}$. As we note in Remark 4.3, this distinction is relevant for full conformal methods, but it holds trivially for split conformal methods due to a separate dataset being used for training and calibration (the symmetry needed is for how the score function treats calibration and test data).
For a more precise distinction, let $\mathcal{S}$ be a real-valued function that takes as its input a point $(x,y)$ and a sequence of other examples $(z_{1},...,z_{n})$, so that $\mathcal{S}\big{(}(x,y),(z_{1},...,z_{n})\big{)}$ can be interpreted as how “nonconforming” or “strange” $(x,y)$ is relative to the examples. Then, we can define symmetric and nonsymmetric score functions as follows: [^10] Note that these definitions implicitly assume that any randomness is fixed, e.g., so that no equality or inequality is trivial due to differing random seeds.
  • $\mathcal{S}$ is a symmetric score function if $\mathcal{S}\big{(}(x,y),(z_{1},...,z_{n})\big{)}=\mathcal{S}\big{(}(x,y),(z_{\sigma(1)},...,z_{\sigma(n)})\big{)}$ for all $n\geq 1$, all permutations $\sigma$ of the indices $\{1,...,n\}$, all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, and all $(z_{1},...,z_{n})=((x_{1},y_{1}),...,(x_{n},y_{n}))$.
  • Otherwise, $\mathcal{S}$ is a nonsymmetric score function; for example, if $\mathcal{S}$ is sensitive to the ordering of its examples such that $\mathcal{S}\big{(}(x,y),(z_{1},...,z_{n})\big{)}\neq\mathcal{S}\big{(}(x,y),(z_{\sigma(1)},...,z_{\sigma(n)})\big{)}$ for some permutation $\sigma$ of the indices $\{1,...,n\}$.
In other words, symmetric score functions treat their training examples $(z_{1},...,z_{n})$ symmetrically, so they can equivalently be defined (as in Section 2) by requiring the examples to be a bag or multiset (rather than a sequence), wherein there is no ordering to the elements. In contrast, nonsymmetric score functions may be sensitive to the ordering of their training inputs.
The proof of Theorem 4.1 provided in Appendix A.1 relies on $\mathcal{S}$ being a symmetric score function to equate score probabilies to data probabilities; that is, the equation $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}=\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ relies on score-function symmetry. To see this, first recall that we justified this equality by stating that (conditioned on $E_{z}$), the fitted and symmetric score function $\mathcal{\widehat{S}}$ induces a bijection between the data points $z_{1},...,z_{n+1}$ and the scores $v_{1},...,v_{n+1}$. As we defined $\mathcal{\widehat{S}}(x,y)=\mathcal{S}\big{(}(x,y),z_{1:(n+1)}\big{)}$, where $z_{1:(n+1)}=\{z_{1},...,z_{n+1}\}$ (an unordered set), this is equivalent to stating that the unfitted (and symmetric) score function $\mathcal{S}$ induces a bijection between the $n+1$ “data objects” $(z_{1},z_{1:(n+1)}),...,(z_{n+1},z_{1:(n+1)})$ and the $n+1$ scores $v_{1},...,v_{n+1}$.
On the other hand, if we allow $\mathcal{S}$ to be a nonsymmetric score function, then in general it can be the case that $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}\neq\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$. This can be seen for instance by comparing the events $E_{v}$ and $E_{z}$: observing an unordered set of data values (i.e., the event $E_{z}$) could result in many different possible sets of scores from a nonsymmetric score function. For nonsymmetric $\mathcal{S}$, we thus are not able to make this substitution to equate score probabilities $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}$ to data probabilities $\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$.
However, a similar argument to the proof in Appendix A.1 could proceed without making this substitution. The main difference is that, rather than the corresponding general result requiring an arbitrary joint PDF $f$ of the data in Eq. (10), instead a the result would require an arbitrary joint PDF of the score values. By making this observation, we aim to point out that extensions to allow for nonsymmetric score functions could be a interesting direction, but one we leave for future work. Such future work could potentially leverage ideas from the “swap step” in Barber et al. (2023), which the authors introduced to accomodate nonsymmetric algorithms.
Appendix B Deriving Weighted Conformal Validity Guarantees and Algorithms for Data under MFCS
This section describes in detail how one can use the general procedure presented in Section 4.3 to derive conformal prediction validity guarantees and practical algorithms for multistep feedback covariate shift (MFCS).
B.1 Deriving MFCS Conformal Validity Guarantees (Sketch Given in Section 4.3)
Here we use the general procedure outlined in Section 4.3 to derive conformal validity guarantees for MFCS.
Step 1: List assumptions
We begin by restating the formal assumptions we used to characterize MFCS:
[…]
$Z_{n+t}=(X_{n+t},Y_{n+t})$ $\hskip 1.42262pt\sim P_{X;Z_{1:(n+t-1)}}^{(t)}\times P_{Y\mid X},\;t=1,\dots,T.$
Recall that informally, these assumptions are stating that (after an IID initialization of $n$ datapoints) the distribution of the covariates $X$ at time $t$ can change arbitrarily depending on past observations $Z_{1:(n+t-1)}=\{Z_{1},...,Z_{n+t-1}\}$, while the conditional label distribution $Y\mid X$ is assumed to remain invariant.
Step 2: Factorize $f$
Next, we factorize the joint probability density function (PDF) $f$ into “dynamic” and “invariant” factors—specifically, into factors that are dependent on versus are invariant to permutations of the data indices. Because in MFCS it is useful to interpret the data indices as timesteps (where the $n$ IID initialized points as distinct observations all occuring at timestep $t=0$), we can think of this distinction as one of time-dependence versus time-invariance.
First, for simplicity, let us focus on factorizing out terms for the most recent timestep $n+t$:
$f$ $(z_{1},...,z_{n+t})$
$=\mathbb{P}(Z_{1}=z_{1},...,Z_{n+t}=z_{n+1})$ More explicit notation
$=\mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1},X_{n+t}=x_{n+t},Y_{n+t}=y_{n+t})$ $Z_{n+t}=(X_{n+t},Y_{n+t})$
$\ \begin{matrix}[l]=\mathbb{P}(Y_{n+t}=y_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1},X_{n+t}=x_{n+t})\\ \qquad\cdot\ \mathbb{P}(X_{n+t}=x_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\\ \qquad\cdot\ \mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\end{matrix}$ Bayes or chain rule
$\ \begin{matrix}[l]=\mathbb{P}(Y_{n+t}=y_{n+t}\ |\ X_{n+t}=x_{n+t})\\ \qquad\cdot\ \mathbb{P}(X_{n+t}=x_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\\ \qquad\cdot\ \mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\end{matrix}$ $\begin{matrix}\text{MFCS invariance assumption:}\\ Y\ |\ X\text{ invariant }\implies\\ (Y_{n+t}\ |\ X_{n+t})\perp\!\!\!\perp Z_{1},...,Z_{n+t-1}\end{matrix}$
These steps are not specific to $n+t$, however; we can repeat the same procedure for $\mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})$ and the index $n+t-1$, and so on. When these steps are performed for all indices $i\in\{1,...,n+t\}$, we obtain
$f$ $(z_{1},...,z_{n+t})$
$=\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(X_{j}=x_{j}\ |\ Z_{1}=z_{1},...,Z_{j-1}=z_{j-1})\cdot\mathbb{P}(Y_{j}=y_{j}|X_{j}=x_{j})\Big{]}$
$=\prod_{j=1}^{n+t}\Big{[}\underbrace{\mathbb{P}(X_{j}=x_{j}\ |\ Z_{1}=z_{1},...,Z_{j-1}=z_{j-1})}_{\begin{matrix}\text{Time-dependent factors}\end{matrix}}\Big{]}\cdot\underbrace{\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(Y_{j}=y_{j}|X_{j}=x_{j})\Big{]}}_{\begin{matrix}\text{Time-invariant factor}\end{matrix}},$
that is, we obtain the factorization into time-dependent and time-invariant factors provided in the main paper Eq. (7).
Step 3: Compute or estimate weights
Lastly, we plug the result of our factorization from Step 2 into Eq. (4) to compute or estimate the conformal weights (for the calibration and test points). Recall that with slightly modified notation from our proof in Appendix A.1 to accommodate $n+t$ data indices, we let $E_{z}^{(t)}$ denote the event $\{Z_{1},...,Z_{n+t}\}=\{z_{1},...,z_{n+t}\}$.
$\mathbb{P}$$\{Z_{n+t}=z_{i}\mid E_{z}^{(t)}\}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}f(z_{\sigma(1)},...,z_{\sigma(n+t)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+t)})}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(Y_{j}=y_{\sigma(j)}\ |\ X_{j}=x_{\sigma(j)})\Big{]}}{\sum_{\sigma}\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})\Big{]}\cdot\prod_{j=1}^{n+t}\Big{[}\mathbb{P}(Y_{j}=y_{\sigma(j)}\ |\ X_{j}=x_{\sigma(j)})\Big{]}}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})},$
where this result is equivalent to Eq. (4.3) in the main paper; it follows because $\prod_{j=1}^{n+t}\mathbb{P}(Y_{j}=y_{\sigma(j)}\ |\ X_{j}=x_{\sigma(j)})$ is invariant to permutations $\sigma$, and thus cancels out in the ratio. For more concise notation to refer to these weights derived using MFCS assumptions, let us define
$\mathbb{P}_{n+t}^{(\text{MFCS})}\{z_{i}\mid E_{z}^{(t)}\}:=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod_{j=1}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{j=1}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})}.$ (14)
A weighted conformal prediction algorithm with these weights $\mathbb{P}_{n+t}^{(\text{MFCS})}\{z_{i}\mid E_{z}^{(t)}\}$ for $i\in\{1,...,n+t\}$ has a valid coverage guarantee as a corollary [^11] We call this result a “corollary” rather than a “theorem” primarily to emphasize that it follows from our main result, Theorem 4.1, when assuming data are generated under MFCS; however, we encourage future authors to use the term they find most fitting for this or other conformal validity guarantees derived from Theorem 4.1. to Theorem 4.1, premised on the MFCS assumptions. This is because our derivation only relied on standard probability rules that hold for any joint PDF [^12] Or more generally, for any valid Radon-Nikodym derivative with respect to an arbitrary base measure; this also includes discrete probability mass functions (PMFs) and mixtures of PDFs and PMFs. $f$ along with the MFCS assumptions listed in Step 1. We now state this corollary for completeness.
Corollary B.1.
Assume that $Z_{i}=(X_{i},Y_{i})\in\mathbb{R}^{d}\times\mathbb{R},i=1,...,n+t$ have the joint PDF $f$ and are generated under multistep feedback covariate shift as in Eq. (3). For any measurable score function $\mathcal{S}$, and any $\alpha\in(0,1)$, define the generalized conformal prediction set (based on $n$ calibration samples) at a point $x\in\mathbb{R}^{d}$ by
$\widehat{\mathcal{C}}_{n}(x)=\bigg{\{}y\in\mathbb{R}:V_{n+t}^{(x,y)}\leq Q_{1-\alpha}\bigg{(}\sum_{i=1}^{n}\mathbb{P}_{n+t}^{(\text{MFCS)}}\{Z_{i}\mid E_{z}\}\cdot\delta_{V_{i}^{(x,y)}}+\mathbb{P}_{n+t}^{(\text{MFCS)}}\{Z_{n+t}\mid E_{z}\}\cdot\delta_{\infty}\bigg{)}\bigg{\}}$ (15)
where $V_{i}^{(x,y)},i\in\{1,...,n+t\}$ are as in (5) and $\mathbb{P}_{n+t}^{(\text{MFCS)}}\{Z_{i}\mid E_{z}\}$ is defined in (14). Then, $\widehat{\mathcal{C}}_{n}$ satisfies
$\mathbb{P}\big{\{}Y_{n+1}\in\widehat{\mathcal{C}}_{n}(X_{n+1})\big{\}}\geq 1-\alpha.$
B.2 Deriving Practical Conformal Algorithms under Agent-Induced MFCS
With the conformal coverage validity guarantee for MFCS in Corollary B.1 at hand, we now turn to deriving practical algorithms for computing or estimating the conformal prediction set defined in Eq. (15) under ML-agent-induced MFCS. As we discuss in Section 4.3 of our main paper, the primary practical bottleneck in this case is computational complexity: While the epistemic challenge is overcome when MFCS is agent-induced (due to $\mathbb{P}(X_{j}\mid Z_{1},...,Z_{j-1})$ representing an ML-agent-controlled query function at time $j$), the complexity for computing the weights in Eq. (14) (equivalent to our main paper’s Eq. (4.3)) is still $\mathcal{O}(\prod_{j=1}^{t}(n+j))$, which quickly becomes intractable for large $t$. To alleviate this bottleneck, we thus proposed an estimation that only uses the “highest-order” terms from the $d$ most recent timesteps,
$\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t+1-d}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t+1-d}}^{n+t}\mathbb{P}(X_{j}=x_{\sigma(j)}\ |\ Z_{1}=z_{\sigma(1)},...,Z_{j-1}=z_{\sigma(j-1)})},$
where we call $d$ our “estimation depth” and $\widehat{p}_{n+t}^{\ (d)}\{z_{i}\mid E_{z}^{(t)}\}$ is our “$d$-step” approximation of the MFCS weights. This $d$-step approximation of the MFCS thus has reduced complexity $\mathcal{O}(\prod_{j=t+1-d}^{t}(n+j))$, which is a tractable polynomial when $d$ is small.
When the ML model controlling the query function $\mathbb{P}(X_{j}\mid Z_{1},...,Z_{j-1})$ treats its training data $Z_{1},...,Z_{j-1}$ symmetrically at each timestep (as in our experiments), we can factorize this even further and derive a recursive algorithm for computing the weights, wherein $d$ corresponds to the recursion depth. To provide intuition for this derivation and recursive implementation, let us start by letting $d=3$ and return to the simpler notation used in the main text, which drops the random variables (i.e., leaves them implicit to be more concise):
$\widehat{p}_{n+t}^{\ ({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3})}\{z_{i}\mid E_{z}^{(t)}\}=\frac{\sum_{\sigma:\sigma(n+t)=i}\prod^{n+t}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t-2}}\mathbb{P}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}{\sum_{\sigma}\prod_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}j=n+t-2}}^{n+t}\mathbb{P}(x_{\sigma(j)}\ |\ z_{\sigma(1)},...,z_{\sigma(j-1)})}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{\sigma}\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
where the second line simply writes out the product. Next, it will soon help us to write the denominator’s summation $\sum_{\sigma}$ equivalently as $\sum_{i}\sum_{\sigma:\sigma(n+t)=i}$, for $i\in\{1,...,n+t\}=[n+t]$
$\widehat{p}_{n+t}^{\ (3)}\{z_{i}\mid E_{z}^{(t)}\}$
$=\frac{\sum_{\sigma:\sigma(n+t)=i}\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{i}\sum_{\sigma:\sigma(n+t)=i}}\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
as we will now make use of knowing that $\sigma(n+t)=i$ within each of the summations $\sum_{\sigma:\sigma(n+t)=i}$ in the numerator and denominator. To do so, we need slightly more notation: for any subset of the data indices $K\subseteq\{1,...,n+t\}=[n+t]$, we write $z_{-K}=\{z_{i}:i\in[n+t]\backslash K\}$ to denote the set of data whose indices are not in $K$; for example, $z_{-\{i,j\}}=\{1,...,n+t\}\backslash\{i,j\}$. And, since we are assuming ML models that treat their data symmetrically at each timestep, we allow our conditional probabilities to ignore the ordering of data observations being conditioned on (e.g., $\mathbb{P}(x\mid z_{-K})$ does not depend on the ordering of the indices in $z_{-K}$). So, with this assumption and notation, the fact that $\sigma(n+t)=i$ within the summation $\sum_{\sigma:\sigma(n+t)=i}$ implies $\mathbb{P}(x_{\sigma(n+t)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-1)})=\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})$. So, we can make this substitution and factor out $\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})$ from $\sum_{\sigma:\sigma(n+t)=i}$, since it no longer depends on $\sigma$:
$\widehat{p}_{n+t}^{\ ({3})}$$\{z_{i}\mid E_{z}^{(t)}\}$
$=$ $\frac{\sum_{\sigma:\sigma(n+t)=i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})}\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\sum_{\sigma:\sigma(n+t)=i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})}\cdot\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})}\sum_{\sigma:\sigma(n+t)=i}\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})}\sum_{\sigma:\sigma(n+t)=i}\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}.$
Next, observe that the index $\sigma(n+t)$ no longer appears within the summation $\sum_{\sigma:\sigma(n+t)=i}$ in either the numerator or denominator (it has been factored out). So, similarly as we did before in the denominator, we can equivalently write the summation $\sum_{\sigma:\sigma(n+t)=i}$ equivalently as $\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}$, for
$\widehat{p}_{n+t}^{\ ({3})}$$\{z_{i}\mid E_{z}^{(t)}\}$
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}}\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}}\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})},$
and similarly as before, within the summation $\sigma(n+t-1)=j$, so $\mathbb{P}(x_{\sigma(n+t-1)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-2)})=\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})$. Making this substitution and factoring, and then repeating the pattern again for the third term ultimately gives
$\widehat{p}_{n+t}^{\ ({3})}$$\{z_{i}\mid E_{z}^{(t)}\}$
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}\sum_{\sigma:\sigma(n+t-1)=j}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{\sigma:\sigma(n+t-1)=j}\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{\sigma:\sigma(n+t-1)=j}\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{k\neq i,j}\sum_{\sigma:\sigma(n+t-2)=k}}\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{k\neq i,j}\sum_{\sigma:\sigma(n+t-2)=k}}\mathbb{P}(x_{\sigma(n+t-2)}\ |\ z_{\sigma(1)},...,z_{\sigma(n+t-3)})}$
$=$ $\frac{\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{k\neq i,j}\mathbb{P}(x_{k}\ |\ z_{-\{i,j,k\}})}{\sum_{i}\mathbb{P}({x_{i}}\ |\ {z_{-\{i\}}})\sum_{j\neq i}{\mathbb{P}(x_{j}\ |\ z_{-\{i,j\}})}\cdot\sum_{k\neq i,j}\mathbb{P}(x_{k}\ |\ z_{-\{i,j,k\}})}.$
More generally, for an arbitrary estimation depth $d\in[n+t]$ we have the following, with annotations added to indicate how increasing our the depth of recursion $d$ will influence the computation,
$\widehat{p}_{n+t}^{\ ({d})}\{z_{i_{1}}\mid E_{z}^{(t)}\}=\frac{\overbrace{\mathbb{P}({x_{i_{1}}}\ |\ {z_{-\{i_{1}\}}})}^{\text{recursion depth 1}}\overbrace{\textstyle\sum_{i_{2}\neq i_{1}}{\mathbb{P}(x_{i_{2}}\ |\ z_{-\{i_{1},i_{2}\}})}}^{\text{recursion depth 2}}\cdots\overbrace{\textstyle\sum_{i_{d}\neq i_{1},...,i_{d-1}}\mathbb{P}(x_{i_{d}}\ |\ z_{-\{i_{1},i_{2},...,i_{d}\}})}^{\text{recursion depth }d}}{\sum_{i_{1}}\mathbb{P}({x_{i_{1}}}\ |\ {z_{-\{i_{1}\}}})\sum_{i_{2}\neq i_{1}}{\mathbb{P}(x_{i_{2}}\ |\ z_{-\{i_{1},i_{2}\}})}\cdots\sum_{i_{d}\neq i_{1},...,i_{d-1}}\mathbb{P}(x_{i_{d}}\ |\ z_{-\{i_{1},i_{2},...,i_{d}\}})},$
and where the denominator can be obtained by summing over all the numerator values for $i_{1}\in[n+t]$.
Appendix C Details on the Roles of Exchangeability, Weighted Exchangeability, and Pseudo-Exchangeability in Conformal Prediction
C.1 The Role of Exchangeability in Conformal Prediction
Summary: Standard conformal prediction can be viewed as an inverted permutation test, where exchangeability allows us to “weigh” or count each permutation equally.
Appendix A.6 of Tibshirani et al. (2019) provides an alternate proof to a key Lemma underlying (standard) conformal prediction’s validity, which essentially presents conformal prediction as an inverted permutation test (though they do not explicitly state this). We reproduce most of that argument here, with some additional explanatory commentary to highlight the role played by the exchangeability assumption in standard conformal prediction.
The general strategy in Tibshirani et al. (2019) is “to condition on the unlabeled multiset of values $\{v_{1},...,v_{n+1}\}$ obtained by our random variables $V_{1},...,V_{n+1}$, and then inspect the probabilities that the last random variable $V_{n+1}$ attains each one of these values.” That is, Tibshirani et al. (2019) condition on the event $\{V_{1},...,V_{n+1}\}=\{v_{1},...,v_{n+1}\}$, which they denote by $E_{v}$—importantly, this is not saying that $V_{i}=v_{i}$ for all $i\in\{1,...,n+1\}$. Rather, it is only saying that the set or “bag” [^13] For simplicity, Tibshirani et al. (2019) assume there are almost surely no ties among those scores (to work with sets rather than multisets/bags); the argument applies in the general case but with more cumbersome notation. of values $\{v_{1},...,v_{n+1}\}$ are the values obtained by the random variables $V_{1},...,V_{n+1}$; we do not yet know whether $v_{1}$ is the value obtained by $V_{1}$, or by $V_{2}$, or by $V_{n+1}$, etc.
With this notation, Tibshirani et al. (2019) consider the following probability:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\},i=1,...,n+1,$
that is, the probability that the value obtained by $V_{n+1}$ is $v_{i}$, conditioned on $E_{v}$ (conditioned on the bag of values).
Next is where Tibshirani et al. (2019) implicitly use an inverted permutation test. With $f$ denoting the joint density function, they state the following, which is analogous to Eq. (4) in our main paper:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}=\frac{\sum_{\sigma:\sigma(n+1)=i}f(v_{\sigma(1)},...,v_{\sigma(n+1)})}{\sum_{\sigma}f(v_{\sigma(1)},...,v_{\sigma(n+1)})}.$
This statement can be thought of as an inverted permutation test as it equates the probability $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}$ to the total probability mass of all permutations $\sigma$ that map $n+1$ to $i$, normalized across all possible permutations of the $n+1$ values. A more explicit way of writing this could include the random variables in the joint density function’s notation, to emphasize that the permutation is on the values and not on the random variables:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}=\frac{\sum_{\sigma:\sigma(n+1)=i}f(V_{1}=v_{\sigma(1)},...,V_{n+1}=v_{\sigma(n+1)})}{\sum_{\sigma}f(V_{1}=v_{\sigma(1)},...,V_{n+1}=v_{\sigma(n+1)})}.$
Note that each argument in $f$ still corresponds to the index of the random variable (not the value); meanwhile, the positions of the observed values are permuted to argument locations.
Next, the key idea is then that—due to exchangeability—every permutation of the data is equally likely, and so the probability of the event $\{V_{n+1}=v_{i}\ |\ E_{v}\}$ can be computed as the total number of permutations (of the $n+1$ values) where $\sigma(n+1)=i$, divided by the total number of all possible permutations (of the $n+1$ values). I.e., by exchangeability $f(V_{1}=v_{\sigma(1)},...,V_{n+1}=v_{\sigma(n+1)})=f(V_{1}=v_{1},...,V_{n+1}=v_{n+1})$ for all $\sigma$, so we can simplify to reduce our statement to counting permutations:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{v}\}$ $=\frac{\sum_{\sigma:\sigma(n+1)=i}f(V_{1}=v_{1},...,V_{n+1}=v_{n+1})}{\sum_{\sigma}f(V_{1}=v_{1},...,V_{n+1}=v_{n+1})}$
$=\frac{\sum_{\sigma:\sigma(n+1)=i}1}{\sum_{\sigma}1}$
$=\frac{n!}{(n+1)!}=\frac{1}{n+1}.$
That is, the third line follows from the second because (in the denominator’s summation) there are $(n+1)!$ possible permutations of $n+1$ values and (in the numerator’s summation) there are $n!$ possible permutations of $n$ values (since in the numerator’s summation we only consider $\sigma$ where $\sigma(n+1)=1$, there are only $n$ values unknown).
Tibshirani et al. (2019) then explain that this result shows that the distribution of $V_{n+1}|E_{z}$ is uniform on the set $\{v_{1},...,v_{n+1}\}$, that is,
$V_{n+1}|E_{v}\sim\frac{1}{n+1}\sum_{i=1}^{n+1}\delta_{v_{i}},$
where $\delta_{v}$ denotes a point mass at the value $v$. The remaining steps of the proof then follow by the definition of a quantile, by the meaning of conditioning on $E_{z}$, and by marginalizing.
The role of exchangeability is to allow us to count every permutation equally in the inverted permutation test. In the next sections, we will see that under weighted exchangeability, we instead need to use the weight functions to see how much “importance” to place on each of the permutations.
C.2 The Role of Weighted Exchangeability in Conformal Prediction
Summary: Weighted CP can be viewed as an inversion of weighted permutation tests, where the weight functions tell us how much to “weigh” or count each permutation.
First, we restate the definition of weighted exchangeability from Tibshirani et al. (2019):
Definition 1: Random variables $V_{1},...,V_{n}$ are said to be weighted exchangeable, with weight functions $w_{1},...,w_{n}$, if the density [^14] Tibshirani et al. (2019) give the following footnote here: “As before, $f$ may be the Radon-Nikodym derivative with respect to an arbitrary base measure.” $f$, of their joint distribution can be factorized as
$f(v_{1},...,v_{n})=\prod_{i=1}^{n}w_{i}(v_{i})\cdot g(v_{1},...,v_{n})$ (16)
where $g$ is any function that does not depend on the ordering of its inputs, i.e., $g(v_{\sigma(1)},...,v_{\sigma(n)})=g(v_{1},...,v_{n})$ for any permutation $\sigma$ of $1,...,n$.
This section sketches the argument in appendix A.7 of Tibshirani et al. (2019), which proves a key lemma underlying the validity of weighted CP, with the goal of understanding the role of the weight functions in an inversion of a weighted permutation test.
Similarly as before and as in equation (4), [^15] Again, for simplicity, Tibshirani et al. (2019) assume that there are almost surely no ties in the values $\{v_{1},...,v_{n+1}\}$ to work with sets rather than multisets/bags, but the results holds more generally (with more cumbersome notation).
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=$$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$
$=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}.$
Assuming that the data are weighted exchangeability, this becomes
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{\sigma(1)},...,z_{\sigma(n+1)})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})g(z_{\sigma(1)},...,z_{\sigma(n+1)})}.$
We can then write out this statement more explicitly, as we did before, by including the random variables in the weight and factor function arguments:
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})g(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})g(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})}.$
In this more explicit version of the statement, we see that $w_{j}$ (for each $j\in\{1,...,n+1\}$) does not only take as its argument the value $z_{\sigma(j)}$, but rather it is a function of the event $Z_{j}=z_{\sigma(j)}$. (The reliance of $w_{j}$ on $Z_{j}$ in the original notation is implicit in the index $j$.) Our more explicit notation can thus allow us to see that—so long as the data are weighted exchangeable—the scalar quantity $\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})$ can be thought of as how much more likely it would be to observe $Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)}$ (given $E_{z}$), relative to the measure that our core function $g$ assigns to this event—i.e., relative to $g(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})$. Put more simply, $\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})$ tells us how much we should “weigh” or “count” a given permutation $\sigma$.
Simplifying using the fact that $g(Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)})=g(Z_{1}=z_{1},...,Z_{n+1}=z_{n+1})$ under weighted exchangeability yields
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})g(Z_{1}=z_{1},...,Z_{n+1}=z_{n+1})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})g(Z_{1}=z_{1},...,Z_{n+1}=z_{n+1})}$
$=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})}{\sum_{\sigma}\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})}.$
This last line shows explicitly that $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}$ is given by the total “weight” of the permutations $\sigma$ such that $\sigma(n+1)=i$, normalized over the total weight of all possible permutations. (We stop here because our emphasis is on the role of the weight functions; the rest of the argument follows similarly as in the previous subsection.)
To review: For each $j\in\{1,...,n+1\}$, the weight function quantity $w_{j}(z_{\sigma(j)})=w_{j}(Z_{j}=z_{\sigma(j)})$ weighs the relative likelihood or “importance” of the event $Z_{j}=z_{\sigma(j)}$, while the product $\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})$ weighs the the relative likelihood or “importance” of the joint event of observing all the data in a particular ordering or permutation $\sigma$ (given $E_{z}$, the observed set of data values), that is of observing $Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)}$. So, the product $\prod_{j=1}^{n+1}w_{j}(Z_{j}=z_{\sigma(j)})$ thus tells us how much to “count” a given permutation $\sigma$.
C.3 The Role of Pseudo-Exchangeability in Conformal Prediction
Summary: Pseudo-exchangeability, as defined in Fannjiang et al. (2022), has a similar role as weighted exchangeability but with “factor functions” in the place of weight functions, and where the factor functions are written to explicitly depend on other datapoints. While pseudo-exchangeability describes one-step feedback covariate shift as a special case (as it was seemingly introduced mainly for this purpose), it does not formally describe MFCS, as an artifact of the definition.
First, we recall the definition of pseudo-exchangeability from (the appendix of) Fannjiang et al. (2022):
Definition 1: Random variables $V_{1},...,V_{n+1}$ are pseudo-exchangeable with factor functions $g_{1},...,g_{n+1}$ and core function $h$ if the density, $f$, of their joint distribution can be factorized as
$f(v_{1},...,v_{n+1})=\prod_{i=1}^{n+1}g_{i}(v_{i};v_{-i})\cdot h(v_{1},...,v_{n+1})$ (17)
where $v_{-i}=v_{1:(n+1)}\backslash v_{i}$, each $g_{i}(\cdot;v_{-i})$ is a function that depends on the multiset $v_{-i}$ (that is, on the values in $v_{-i}$ but not on their ordering), and $h$ is a function that does not depend on the ordering of its $n+1$ inputs.
Note that pseudo-exchangeability is similar to weighted exchangeability, except the “factor functions” $g_{i}(v_{i};v_{-i})$ take the place of the weight functions $w_{i}(v_{i})$ to explicitly describe how $v_{i}$ can depend (but only symmetrically, as per the definition) on the other observations $v_{-i}$. The analogous argument in Fannjiang et al. (2022) is very similar to that of Tibshirani et al. (2019) except using pseudo-exchangeability. [^16] Again, for simplicity, Fannjiang et al. (2022) assume that there are almost surely no ties to work with sets rather than multisets/bags, but the results holds more generally (with more cumbersome notation). As before, we begin with the statement of equation (4):
$\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=$$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$
$=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}f(z_{\sigma(1)},...,z_{\sigma(n+1)})}{\sum_{\sigma}f(z_{\sigma(1)},...,z_{\sigma(n+1)})},$
Using pseudo-exchangeability, this becomes
$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})h(z_{\sigma(1)},...,z_{\sigma(n+1)})}{\sum_{\sigma}\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})h(z_{\sigma(1)},...,z_{\sigma(n+1)})},$
where the core function $h$ does not depend on the ordering of its inputs, so we have
$\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})h(z_{1},...,z_{n+1})}{\sum_{\sigma}\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})h(z_{1},...,z_{n+1})}$
$=$ $\frac{\sum_{\sigma:\sigma(n+1)=i}\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})}{\sum_{\sigma}\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})}.$ (18)
Here, in the same way that for weighted exchangeable data the product $\prod_{j=1}^{n+1}w_{j}(z_{\sigma(j)})$ represents the (relative) likelihood of observing the joint sequence of events $Z_{1}=z_{\sigma(1)},...,Z_{n+1}=z_{\sigma(n+1)}$, for pseudo-exchangeable data the analogous role is played by the product $\prod_{j=1}^{n+1}g_{j}(z_{\sigma(j)};z_{-\sigma(j)})$: the role of this product is to tell us how much to “weigh” or “count” a given permutation $\sigma$. However, the specification in pseudo-exchangeability’s definition that “$g_{i}(\cdot;v_{-i})$ is a function that depends on the multiset $v_{-i}$ (that is, on the values in $v_{-i}$ but not on their ordering)” (emphasis added) prevents pseudo-exchangeability from formally describing multistep feedback covariate shift. We state this more formally in the following remark.
Remark C.1.
Pseudo-exchangeability does not formally characterize multistep FCS. Consider a two-step instance of FCS for a counterexample, that is with data $Z_{1},...,Z_{n+2}$ generated under MFCS. The factor function for the first step $g_{n+1}(X_{n+1}=x_{n+1}\ ;\ Z_{1}=z_{1},...,Z_{n}=z_{n})$ depends on $Z_{1},...,Z_{n}$ differently than it does on $Z_{n+2}$ (i.e., it does not depend on $Z_{n+2}$ at all). So, the factor function $g_{n+1}(\cdot\ ;\ Z_{1}=z_{1},...,Z_{n}=z_{n})$ cannot be said to depend on the multiset of values $z_{-(n+1)}=\{z_{1},...,z_{n},z_{n+2}\}$ independently of the ordering of the values (i.e., it requires knowing which of the values corresponds to $Z_{1:n}$ versus $Z_{n+2}$), as is required by the defintion of pseudo-exchangeability from Fannjiang et al. (2022).
Appendix D Additonal Experimental Details
D.1 Black-box Optimization Experimental Details
Protein Design Datasets Our multistep biomolecular design experiments leverage fluorescent protein data from Poelwijk et al. (2019), previously used for evaluation of related single-step FCS CP methods in Fannjiang et al. (2022) and Prinster et al. (2023). This dataset has the benefit of a combinatorially complete set of labels, which simulates measuring a design in an experiment. In particular, both a “blue” and “red” wavelength fluorescence strength were experimentally measured for each of $1922^{13}=8,$ 192 possible combinations of binary variations to a wild-type (natural) fluorescent protein at 13 positions on the primary sequence. As in Fannjiang et al. (2022), we set the noise scale for each sequence $i$ to the residual norm between that sequence’s label and a separately fit linear model.
D.2 Active Learning Experimental Details
Datasets for Active Learning Experiments We evaluate on four datasets from the UCI Machine Learning Repository (Frank, 2010) that are commonly used for evaluation in the CP literature (Tibshirani et al., 2019; Barber et al., 2021; Prinster et al., 2022, 2023): the NASA airfoil self-noise dataset (1503 samples, $p=5$) (Brooks et al., 2014), the communities and crime dataset (1994 samples, $p=99$) (Redmond, 2009), the Medical Expenditure Panel Survey 2016 data set [^17] https://meps.ahrq.gov/mepsweb/data_stats/download_data_files_detail.jsp?cboPufNumber=HC-192 (33005 samples, $p=107$) (preprocessed as in Barber et al. (2021); details for an older version of the data are in Ezzati-Rice et al. (2008)), and the blog feedback dataset (60021 samples, $p=281$) (Buza, 2014).
Model for Active Learning Experiments: The model we used for the active learning experiments is scikit-learn’s GaussianProcessRegressor, with kernel=DotProduct()+WhiteKernel() and default parameters otherwise.
Appendix E Additional Experimental Results
E.1 Bounding the Query Function to Enforce Informative (Finite) Intervals
Because noninformative prediction sets ($\widehat{\mathcal{C}}_{n}(x)=\mathcal{Y}$) occur in weighted CP when a test point’s normalized weight as in Eq. (9) exceeds $\alpha$, we now consider redesigning the query probability functions to prevent this from happening. For simplicity, we impose this constraint on the one-step weights. In particular, we consider bounding the query probability function $\mathbb{P}(x\mid Z_{1},...,Z_{j-1})$ so that instead of being proportional to $\exp(\lambda\cdot u_{t}(x))$, the probabilities are instead proportional to a bounded function:
$\mathbb{P}(x\mid Z_{1},...,Z_{j-1})\propto$ $\min\big{(}\exp(\lambda\cdot u_{t}(x)),B\big{)},$
where $\text{where }B=\sup\bigg{\{}b>0:\frac{b}{\sum_{i=1}^{n}\min\big{(}\exp(\lambda\cdot u_{t}(x)(X_{i})),b\big{)}+b}<\alpha\bigg{\}}.$
That is, we select $B$ as the largest bound that we could use to avoid infinite prediction intervals (in the depth $d=1$ weights). This is because
$\min\big{(}\exp(\lambda\cdot u_{t}(x)),B\big{)}\leq B$
$\implies\frac{\min\big{(}\exp(\lambda\cdot u_{t}(x)),B\big{)}}{\sum_{i=1}^{n}\min\big{(}\exp(\lambda\cdot u_{t}(X_{i})),B\big{)}+\min\big{(}\exp(\lambda\cdot u_{t}(x)),B\big{)}}\leq\frac{B}{\sum_{i=1}^{n}\min\big{(}\exp(\lambda\cdot u_{t}(X_{i})),B\big{)}+B}<\alpha.$
Across the four real-world datasets and the same experimental conditions we used for our active learning experiments in the main paper Figure 3, the following results empirically validate that imposing the proposed adaptive bound on our MFCS CP method’s query function completely prevents any noninformative (infinite width) intervals, all without sacrificing coverage even over a long time horizon (70 active learning steps). For comparison, for each dataset we provide the corresponding results for the original unbounded query function. We moreover plot the “relative magnitude” of the bound $B$—i.e., we plot $B/B_{\max}$, where $B_{\max}=\max_{x\in\text{pool}}(\exp(\lambda\cdot\hat{\sigma}(x)))$—across the same trajectory of active learning steps.
Airfoil dataset
Figure 4: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the airfoil dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.
Communities dataset
Figure 5: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the communities dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.
MEPS dataset
Figure 6: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the MEPS dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.
Blog dataset
Figure 7: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the blog dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.
The bounded query function’s improvement in prediction-interval informativeness sometimes comes at the cost of an initially delayed decrease in MSE—where the constraint forces the agent to initially be more “cautious” in its exploration—though the long-run MSE for the bounded and unbounded query functions are comparable for most datasets. The increasing bound magnitude over time demonstrates that this bounded query function enables the active-learning agent to be less cautious over time by relaxing the constraint of the bound once it has seen more data.